OpenStudy (anonymous):

y varies jointly with p and q. When p = 1 and q = 2, y = 30. Find y when p = 10 and q = 3. (4 points) Answer

OpenStudy (anonymous):

what does "varies jointly" mean? a linear combination?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i don't think that's enough information to solve the problem

OpenStudy (anonymous):

varies jointly is the same as varies directly

OpenStudy (anonymous):

these are the answers y = 60 y = 150 y = 300 y = 450

OpenStudy (anonymous):

go get em! klown!

OpenStudy (anonymous):

y=kpq k=y/(pq)=30/2=15 so y=15pq y=15*10*3=450

OpenStudy (anonymous):

ohh, is that what they mean by varies jointly? y=kpq? I thought they might have meant y = ap + bq in which case there was insufficient info. nice one ghass